Palmer's average running speed is 3 Kilometers per hour faster than his walking speed. If Palmer can run around a 20-Kilometer course in 2 hours, how many hourse would it take her Palmer to walk the same course?
I NEED TO SEE THE WORK ON HOW YOU GOT THE ANSWER!!

Answer by mananth(16946)

walking speed =x
running speed = x+3
d=rt
20=(x+3)*2
20=2x+6
20-6=2x
14=2x
x=7 kmph running speed
so walking speed = 4 kmph (7-3)
d=rt
d=20
r=4
20=4*t
20/4=t
t= 5 hours walking time for 20 km